Need help solving:
sin^2(x/2)-2 = 0 and
sec(x/2) = cos(x/2)
for exact solution over the interval [0,2pi].
Thx!

Answer by stanbon(75887)

sin^2(x/2)-2 = 0
sin^2(x/2) = 2
Ans: no solution since sin^2 cannot exceed +1.
-----------------------------------
sec(x/2) = cos(x/2)
1/cos(x/2) = cos(x/2)
cos^2(x/2) = 1
x/2 = 0 or x/2 = 2pi
x = 0 or x = 4pi
-------------------------
Cheers,
Stan H.
-------------------------